5. Find a general solution to the differential equation using the method of variation of parameters y + 10y + 25y 5e-50
Answer #1

1e lot W(%,%) Now particular fol il y today = -2,64) (hide fift? Itt +24) (3, 40 ft), Wlay, ja where f(t) = 5e-st Wlad) • Now2Ip (H) = { post That general el of diff. equation in 2-57 şte 54 y = (G+G+) est3